# Enhancing Event Extraction from Short Stories through Contextualized Prompts

### v1

@@ Abstract
idelines to annotate the short story dataset. Later, we apply the baseline methods for automatically detecting and categorizing events. We also propose a prompt-based method for event detection and classification. The proposed method outperforms the baselines, while having significant improvement of more than 4% for the class CONFLICT in event classification task. The dataset and the code are publicly available on Github [^1] https://github.com/chaitanyakirti/vrittanta.
@@ 1 Introduction
The process of extracting structured information from a text, such as entities, relations, events, etc., is known as information extraction (IE). Event extraction, a crucial component of IE, is responsible for locating trigger mentions of a specific event type and those events’ participants. An event is a particular occurrence at a particular time and place, involves one or more individuals, and is often defined as a change of states [eventdefACL]. The task of event extraction comprises three subtasks: (1) Event detection (ED) that pulls trigger words from the text to indicate when an event occurs. Trigger words are those words that most clearly express the occurrence of an event; (2) Assigning a predefined event type to event triggers; and (3) Event argument extraction, i.e., to identify an event’s participants. The task is challenging because there are many ways to convey the same event, and same phrase can express distinct events. For example, left could mean a movement, as in Figure 1, or could mean the act of not covering the food, as in “the food has been left uncovered”.
@@
Event extraction is useful for information retrieval [basile2014extending], question answering [yang2003structured], narrative schema extraction [chambers2009unsupervised], reading comprehension and summarization [cheng-erk-2018-implicit], and for building knowledge banks (KBs). Event extraction has several applications in various domains. Structured events, for example, can be used explicitly to extend knowledge bases, allowing for more logical reasoning and inference[eearea1]. Governments have long focused their public affairs management on event identification and tracking, as understanding the outbursts and evolution of popular social events allows them to respond quickly [eearea2]. In the financial domain, event extraction can assist businesses in rapidly identifying consumer reactions to their goods and deriving indications for risk analysis and trading recommendations [eearea3]. Event extraction can be used in the biomedical domain to identify the alterations in biomolecules such as genes and proteins or interactions between two or more biomolecules [eearea4]. In short, advances in event extraction techniques and systems can support a wide range of domains.
@@
Previous works primarily cover datasets in the domain of news, bio-medical etc., involving neural methods for event extraction [sims2019literary]. Event extraction from short stories and the classification of those events remains a less explored area. This motivated us to create a dataset of Indian short stories. Our objective is to make a valuable contribution to resource building and support future research in this domain through the introduction of the Vrittanta-en corpus.
@@
As prompt-based generative methodologies have recently gained prominence across all NLP tasks including for the event extraction, we thoroughly assess the quality of Vrittanta through both (1) neural approaches, and (2) the prompt-based methods. We adhered to prior research in prompt-based generative event extraction models utilizing PLMs with an encoder-decoder transformer architecture, similar to models like BART [lewis-etal-2020-bart], and T5 [JMLR:v21:20-074]. Our proposed methodology employs the BART model as its PLM, which is a conventional Transformer-based PLM featuring both an Encoder and a Decoder. These evaluations provide a comprehensive understanding of the dataset’s reliability and effectiveness, enabling us to validate its suitability for a wide range of applications and research endeavours. Detailed annotation guidelines, annotated dataset, and top-performing models are accessible on GitHub.
@@
The importance of gold-standard data in building Natural Language Processing (NLP) applications cannot be overstated. As NLP strives to handle more intricate linguistic phenomena and diverse labelling tasks, the demand for substantial labelled datasets remains high [sun2017revisiting, human-annotation]. However, the annotation process has become more challenging and costly due to these complexities. Manually annotating linguistic resources for training and testing, particularly in specialized text domains, is a resource-intensive and expensive endeavour. So, to expand the dataset, the best-performing model was employed, whose quality was established experimentally.
@@
The rest of the paper is organized as follows. Section 2 presents previous work. Section 3 presents the dataset construction procedure, followed by the model details in Section 4. Section 5 describes the results and discussions. Finally, Section 7 concludes the paper.
@@ 2 Previous Work
Several public evaluation programs (PEPs) have been initiated to enhance the competitiveness of event extraction, including MUC [sundheim1993survey], TDT [allan2002topic], ACE [ace], TAC-KBP [mcnamee2010evaluation], and SemEval [^2] https://aclanthology.org/venues/semeval/. They have provided standardized evaluation resources, encouraged competition among researchers, and set benchmarks for performance measurement.
@@
The primary requirement of any event extraction task is detecting specific types of events mentioned in the annotated data. In the existing literature, the event extraction task has been divided into open and closed domains [xiang2019survey]. Closed-domain event extraction uses a predefined event schema to discover and extract desired events of a particular type from text [hogenboom2011overview]. ACE 2005 [doddington2004automatic], ERE [song2015light] and TAC-KBP mitamura2015overview use closed domain method by defining the structure and type of events first and then extract results based on the defined structure. The standard dataset ACE 2005 corpus includes 8 types of events, with 33 subtypes [ace].
@@
The lack of annotated corpus of biological literature led to the development of the GENIA corpus [kim2003genia]. It allows NLP techniques to work for text mining at the biomolecular level by providing semantically annotated reference scientific articles. Another corpus in the clinical domain for tracking the disease condition, observing treatment outcomes and complications, and detecting side effects in medication was introduced to the research community by the organizers of the i2b2 Challenge [sun2013evaluating]. Another corpus builds on top of TimeBank, adding to it semantic information for the annotation dealing with the factuality of events in text [sauri2009factbank]. A corpus of sentences is annotated as general or specific by annotators on Amazon Mechanical Turk from news articles comprising reports on events, finance, and science journalism [louis2012corpus]. DBpedia [knuth2015dbpedia] is another dataset created based on live extraction of Wikipedia. The authors follow rule sets to detect events, such as news related to sports, politics, real life, etc., from any update in Wikipedia content. The recognition and elaboration of historical events is a crucial step when dealing with historical texts. A historical corpus named Histo Corpus [sprugnoli2019novel] contains historical text from travel narratives and news, published between the late nineteenth and early twentieth century.
@@
The role of events in literary fiction, however, is very different from their role in fact-based reporting of events in the real world, including historical texts [sprugnoli2019novel]. In literature, imaginary causalities exist, which are very difficult to extract in contrast to the hard-coded causalities in fact-based reports [sauri2009factbank]. Novels tend to be much longer than news articles [louis2012corpus] or clinical documents [nedellec2013overview] and tend to have more complex narrative structures. Since event extraction from historical texts, clinical documents, or news has now become possible due to the availability of datasets, the same was not possible for the literature domain. However, recently,  sims2019literary have annotated a new dataset of literary events, Litbank [^3] https://github.com/dbamman/litbank/tree/master/events. This dataset focuses solely on events in literary texts that are depicted as happening actually. It has taken literary texts from novels published in the nineteenth and twentieth centuries. In terms of techniques used, existing closed-domain approaches can be divided into four categories: pattern matching [riloff1993automatically], machine learning [liao2010using], deep learning [zheng2019doc2edag], and prompt-based approaches.
@@
Many pattern-based event extraction systems have been created for many application areas, including biomedical event extraction [cohen2009high], [casillas2011using] and financial event extraction [borsje2010semi], as a result of the AutoSlog system extracting terrorist incidents [riloff1993automatically]. To extract biological events, Casillas et al. [casillas2011using] used Kybots (Knowledge Yielding Robots). To manually generate event patterns, the Kybots system uses a rule-based approach. Cao et al. [cao2018including] proposed that further event patterns can be generated by integrating the ACE corpus with other expert-defined patterns.
@@
In the machine learning approach, a nearest-neighbor learning method is used for recognizing triggers and a maximum entropy learner(s) classifier is adopted for identifying arguments [ahn2006stages]. Event patterns and token attributes can also be combined as input for machine learning classifiers [meyers2001parsing]. In learning deep or complex nonlinear relations, deep learning-based methods can overcome the limits of classic machine learning methods. A CNN has been shown to be effective at capturing the syntactic and semantics of a sentence [kalchbrenner2014convolutional] because it is capable of learning text-hidden features based on continuous and generalized word embeddings. Nguyen and Grishman [liu2017improving] may be the first to build a CNN for event detection, that is, detecting a trigger and the sort of event it causes in a sentence. As the network input, each word is first turned into a real-valued vector representation that is a concatenation of the word embedding, its location embedding, and the entity type embedding. The CNN consists of a convolution layer, a max pooling layer, and a softmax layer from input to output, and outputs the classification result for each word. The attention mechanism has recently become popular in several NLP tasks [liu2017exploiting]. zeng2016convolution recommended that each word’s local contextual representation be learned first with a CNN, then concatenated with the output of another Bi-LSTM to generate the final word representation for classification. The generative adversarial network (GAN) is another important type of hybrid model, which typically comprises of two neural networks competing against each other [hong2018self].
@@
Prompt-based learning methods, separately treated as a single category, model various sequence-to-sequence tasks with the help of pre-trained Encoder-Decoder-based Transformer architectures like BART [lewis-etal-2020-bart], and T5 [JMLR:v21:20-074]. In contrast to the pre-training and fine-tuning paradigm, which requires a lot of resources and data, prompt-based techniques transform the downstream tasks into a form, more aligned with the model’s pre-training tasks. By creating linked prompts with blanks and determining a mapping from certain filled words to projected categories, schick-schutze-2021-exploiting converts a number of classification tasks into cloze activities. They construct related prompts and try to map the blanks, that have been used for constructing the prompts, to predicted categories. li-liang-2021-prefix mainly work on generation tasks and propose a lightweight prefix-tuning approach where they freeze the model parameters and only adjust a sequence of continuous task-specific vectors. The task in BART [lewis-etal-2020-bart] is extended for event argument extraction, where the events are required to be given as the input to predict its arguments [ma2022prompt, zhong2023contextualized, peng2024prompt]. These models focus on event argument extraction, leaving the tasks of event detection and classification unaddressed. Therefore, there is a gap where the subtasks of event extraction and classification can be completed using prompt-based learning methods. To put it differently, there was almost negligible effort to discover a definition of an event that pertains to the domain of short stories. This contrasts with the approach taken in other fields, such as the news or clinical domain. Moreover, the potential of contextualized prompts has not been leveraged. This serves as the primary reason driving the present study.
@@ 3 Dataset Construction
This section covers the source of the short stories, key annotation guidelines, and dataset statistics. Furthermore, detailed statistics about the event classes used during the annotation process are presented. This extends our previous work, which focused solely on event detection with single-word event triggers [kirti]. In this extension, we explore event classification alongside event detection, expanding our scope to include multi-word event expressions.
@@ 3.1 Source of the short stories
We select 25 stories for annotation from each of the above-mentioned eight sources. This rich assortment of stories helps in exploring various writing styles and narrative techniques for a wide range of events. Each event within these stories is accompanied by an event trigger. These event triggers, typically expressed as verbs, although occasionally as nouns and adjectives, are specifically designed to elicit the occurrence of that particular event.
@@ 3.2 Annotation Guideline
This subsection discusses the key annotation guideline. Short stories possess unique characteristics that set them apart from other types of texts. Detecting events within short stories poses a specific challenge, especially considering their target audience of children and the narrative structures employed. In these stories, it is common to encounter animals or even inanimate objects engaging in conversations with one another. The occurrence of events in such stories relies heavily on the narrative, genre, and storytelling style. Considering the prior research findings, we have created fresh annotation guidelines centered on recognizing and categorizing event mentions to address these specific challenges for events in short stories. The issues and rules addressed by the guidelines established in ACE ace, NewsReader eventdefNRG, HISTO Corpus sprugnoli2019novel, and Litbank lit are taken into account. By leveraging insights from prior research, the aim is to enhance the accuracy and effectiveness of event detection within short stories. The proposed guidelines are valuable resources for annotators, as clear instructions and criteria are provided for identifying and labeling events within the context of children’s short stories. Hereafter, bold words represent realis events, and underlined word shows non-real events that seem to be misunderstood as real events.
@@ 3.2.2 Dimensions to capture realis events
The dimensions followed in the Light ERE approach [aguilar2014comparison] are considered to capture realis events and are mentioned below.
@@
      He traveled to Delhi in the last September.
@@ 3.2.3 Event Classes
Every event that has been annotated needs to be categorized to be a particular class. Event categorization is required to know the distribution of different events present in the domain of short stories. As the events in short stories diverge significantly from earlier datasets, this work could not adhere to the categorization outlined in prior literature. The events are classified into seven types so as to cover all the major happenings. The classes are COGNITIVE-MENTAL-STATE (CMS), COMMUNICATION (COM), CONFLICT (CON), GENERAL-ACTIVITY (GA), LIFE-EVENT (LE), MOVEMENT (MOV), and OTHERS (OTH).
@@
Emotional actions like hugging, smiling, laughing, smiling, etc are also tagged in this type.
@@
V. LIFE-EVENT: Life event are mostly inspired by the ACE(2005)[ace]. These subtypes explore the event tags by LIFE-EVENT.
@@
3.3 Evaluation Metrices
We used the identical evaluation criteria as previous studies [wadden2019entity], which presented the Precision (P), Recall (R), and F1 scores for event detection. We defined TP as a true positive, FN as a false negative, and FP as a false positive.
$P=\frac{TP}{TP+FP}\quad$
$R=\frac{TP}{TP+FN}\quad$
$F1=2*\frac{P*R}{P+R}\quad$

### v2

@@ Abstract
idelines to annotate the short story dataset. Later, we apply the baseline methods for automatically detecting and categorizing events. We also propose a prompt-based method for event detection and classification. The proposed method outperforms the baselines, while having significant improvement of more than 4% for the class CONFLICT in event classification task.
@@ 1 Introduction
The process of extracting structured information from a text, such as entities, relations, events, etc., is known as information extraction (IE). Event extraction, a crucial component of IE, is responsible for locating trigger mentions of a specific event type and those events’ participants. An event is a particular occurrence at a particular time and place, involves one or more individuals, and is often defined as a change of states [Peinelt et al., 2020]. The task of event extraction comprises three subtasks: (1) Event detection (ED) that pulls trigger words from the text to indicate when an event occurs. Trigger words are those words that most clearly express the occurrence of an event; (2) Assigning a predefined event type to event triggers; and (3) Event argument extraction, i.e., to identify an event’s participants. The task is challenging because there are many ways to convey the same event, and same phrase can express distinct events. For example, left could mean a movement, as in Figure 1, or could mean the act of not covering the food, as in “the food has been left uncovered”.
@@
Event extraction is useful for information retrieval [Basile et al., 2014], question answering [Yang et al., 2003], narrative schema extraction [Chambers and Jurafsky, 2009], reading comprehension and summarization [Cheng and Erk, 2018], and for building knowledge banks (KBs). Event extraction has several applications in various domains. Structured events, for example, can be used explicitly to extend knowledge bases, allowing for more logical reasoning and inference[Rospocher et al., 2016]. Governments have long focused their public affairs management on event identification and tracking, as understanding the outbursts and evolution of popular social events allows them to respond quickly [Piskorski et al., 2007]. In the financial domain, event extraction can assist businesses in rapidly identifying consumer reactions to their goods and deriving indications for risk analysis and trading recommendations [Nuij et al., 2013]. Event extraction can be used in the biomedical domain to identify the alterations in biomolecules such as genes and proteins or interactions between two or more biomolecules [Vanegas et al., 2015]. In short, advances in event extraction techniques and systems can support a wide range of domains.
@@
Previous works primarily cover datasets in the domain of news, bio-medical etc., involving neural methods for event extraction [Sims et al., 2019b]. Event extraction from short stories and the classification of those events remains a less explored area. This motivated us to create a dataset of Indian short stories. Our objective is to make a valuable contribution to resource building and support future research in this domain through the introduction of the Vrittanta-EN corpus.
@@
As prompt-based generative methodologies have recently gained prominence across all NLP tasks including for the event extraction, we thoroughly assess the quality of Vrittanta-EN through both (1) neural approaches, and (2) the prompt-based methods. We adhered to prior research in prompt-based generative event extraction models utilizing PLMs with an encoder-decoder transformer architecture, similar to models like BART [Lewis et al., 2020], and T5 [Raffel et al., 2020]. Our proposed methodology employs the BART model as its PLM, which is a conventional Transformer-based PLM featuring both an Encoder and a Decoder. These evaluations provide a comprehensive understanding of the dataset’s reliability and effectiveness, enabling us to validate its suitability for a wide range of applications and research endeavours. Detailed annotation guidelines, annotated dataset, and top-performing models are accessible on GitHub.
@@
The importance of gold-standard data in building Natural Language Processing (NLP) applications cannot be overstated. As NLP strives to handle more intricate linguistic phenomena and diverse labelling tasks, the demand for substantial labelled datasets remains high [Sun et al., 2017, Liu, 2023]. However, the annotation process has become more challenging and costly due to these complexities. Manually annotating linguistic resources for training and testing, particularly in specialized text domains, is a resource-intensive and expensive endeavour. So, to expand the dataset, the best-performing model was employed, whose quality was established experimentally.
@@
The rest of the paper is organized as follows. Section 2 presents previous work. Section 3 presents the dataset construction procedure, followed by the description of EVENT PROMPT model in Section 4. 5 presents the experiments conducted. Section 6 describes the results and discussions. The errors are analysed in Section 7. Finally, Section 8 concludes the paper.
@@ 2 Previous Work
Several public evaluation programs (PEPs) have been initiated to enhance the competitiveness of event extraction, including MUC [Sundheim and Chinchor, 1993], TDT [Allan, 2002], ACE [Consortium and others, 2005], TAC-KBP [McNamee et al., 2010], and SemEval [^1] https://aclanthology.org/venues/semeval/. They have provided standardized evaluation resources, encouraged competition among researchers, and set benchmarks for performance measurement.
@@
The primary requirement of any event extraction task is detecting specific types of events mentioned in the annotated data. In the existing literature, the event extraction task has been divided into open and closed domains [Xiang and Wang, 2019]. Closed-domain event extraction uses a predefined event schema to discover and extract desired events of a particular type from text [Hogenboom et al., 2011]. ACE 2005 [Doddington et al., 2004], ERE [Song et al., 2015] and TAC-KBP Mitamura et al. [2015] use closed domain method by defining the structure and type of events first and then extract results based on the defined structure. The standard dataset ACE 2005 corpus includes 8 types of events, with 33 subtypes [Consortium and others, 2005].
@@
The lack of annotated corpus of biological literature led to the development of the GENIA corpus [Kim et al., 2003]. It allows NLP techniques to work for text mining at the biomolecular level by providing semantically annotated reference scientific articles. Another corpus in the clinical domain for tracking the disease condition, observing treatment outcomes and complications, and detecting side effects in medication was introduced to the research community by the organizers of the i2b2 Challenge [Sun et al., 2013]. Another corpus builds on top of TimeBank, adding to it semantic information for the annotation dealing with the factuality of events in text [Saurí and Pustejovsky, 2009]. A corpus of sentences is annotated as general or specific by annotators on Amazon Mechanical Turk from news articles comprising reports on events, finance, and science journalism [Louis and Nenkova, 2012]. DBpedia [Knuth et al., 2015] is another dataset created based on live extraction of Wikipedia. The authors follow rule sets to detect events, such as news related to sports, politics, real life, etc., from any update in Wikipedia content. The recognition and elaboration of historical events is a crucial step when dealing with historical texts. A historical corpus named Histo Corpus [Sprugnoli and Tonelli, 2019] contains historical text from travel narratives and news, published between the late nineteenth and early twentieth century.
@@
The role of events in literary fiction, however, is very different from their role in fact-based reporting of events in the real world, including historical texts [Sprugnoli and Tonelli, 2019]. In literature, imaginary causalities exist, which are very difficult to extract in contrast to the hard-coded causalities in fact-based reports [Saurí and Pustejovsky, 2009]. Novels tend to be much longer than news articles [Louis and Nenkova, 2012] or clinical documents [Nédellec et al., 2013] and tend to have more complex narrative structures. Since event extraction from historical texts, clinical documents, or news has now become possible due to the availability of datasets, the same was not possible for the literature domain. However, recently,  Sims et al. [2019b] have annotated a new dataset of literary events, Litbank [^2] https://github.com/dbamman/litbank/tree/master/events. This dataset focuses solely on events in literary texts that are depicted as happening actually. It has taken literary texts from novels published in the nineteenth and twentieth centuries. In terms of techniques used, existing closed-domain approaches can be divided into four categories: pattern matching [Riloff and others, 1993], machine learning [Liao and Grishman, 2010], deep learning [Zheng et al., 2019], and prompt-based approaches.
@@
Many pattern-based event extraction systems have been created for many application areas, including biomedical event extraction [Cohen et al., 2009], [Casillas et al., 2011] and financial event extraction [Borsje et al., 2010], as a result of the AutoSlog system extracting terrorist incidents [Riloff and others, 1993]. To extract biological events, Casillas et al. [Casillas et al., 2011] used Kybots (Knowledge Yielding Robots). To manually generate event patterns, the Kybots system uses a rule-based approach. Cao et al. [Cao et al., 2018] proposed that further event patterns can be generated by integrating the ACE corpus with other expert-defined patterns.
@@
In the machine learning approach, a nearest-neighbor learning method is used for recognizing triggers and a maximum entropy learner(s) classifier is adopted for identifying arguments [Ahn, 2006]. Event patterns and token attributes can also be combined as input for machine learning classifiers [Meyers et al., 2001]. In learning deep or complex nonlinear relations, deep learning-based methods can overcome the limits of classic machine learning methods. A CNN has been shown to be effective at capturing the syntactic and semantics of a sentence [Kalchbrenner et al., 2014] because it is capable of learning text-hidden features based on continuous and generalized word embeddings. Nguyen and Grishman [Liu et al., 2017a] may be the first to build a CNN for event detection, that is, detecting a trigger and the sort of event it causes in a sentence. As the network input, each word is first turned into a real-valued vector representation that is a concatenation of the word embedding, its location embedding, and the entity type embedding. The CNN consists of a convolution layer, a max pooling layer, and a softmax layer from input to output, and outputs the classification result for each word. The attention mechanism has recently become popular in several NLP tasks [Liu et al., 2017b]. Zeng et al. [2016] recommended that each word’s local contextual representation be learned first with a CNN, then concatenated with the output of another Bi-LSTM to generate the final word representation for classification. The generative adversarial network (GAN) is another important type of hybrid model, which typically comprises of two neural networks competing against each other [Hong et al., 2018].
@@
Prompt-based learning methods, separately treated as a single category, model various sequence-to-sequence tasks with the help of pre-trained Encoder-Decoder-based Transformer architectures like BART [Lewis et al., 2020], and T5 [Raffel et al., 2020]. The prompts help the model understand the specific requirements of the task and improve its performance by providing the model with relevant context for understanding and extracting events from text. These prompts typically consist of a combination of text and placeholders that guide the model on what information to extract and how to structure the output. In contrast to the pre-training and fine-tuning paradigm, which requires a lot of resources and data, prompt-based techniques transform the downstream tasks into a form, more aligned with the model’s pre-training tasks. By creating linked prompts with blanks and determining a mapping from certain filled words to projected categories, Schick and Schütze [2021] converts a number of classification tasks into cloze activities. They construct related prompts and try to map the blanks, that have been used for constructing the prompts, to predicted categories. Li and Liang [2021] mainly work on generation tasks and propose a lightweight prefix-tuning approach where they freeze the model parameters and only adjust a sequence of continuous task-specific vectors. The task in BART [Lewis et al., 2020] is extended for event argument extraction, where the events are required to be given as the input to predict its arguments [Ma et al., 2022, Zhong et al., 2023, Peng et al., 2024]. These models focus on event argument extraction, leaving the tasks of event detection and classification unaddressed. Therefore, there is a gap where the subtasks of event extraction and classification can be completed using prompt-based learning methods. To put it differently, there was almost negligible effort to discover a definition of an event that pertains to the domain of short stories. This contrasts with the approach taken in other fields, such as the news or clinical domain. Moreover, the potential of contextualized prompts has not been leveraged. This serves as the primary reason driving the present study.
@@ 3 Dataset Construction
This section covers the source of the short stories, key annotation guidelines, and dataset statistics. Furthermore, detailed statistics about the event classes used during the annotation process are presented. This extends our previous work, which focused solely on event detection with single-word event triggers [Kirti et al., 2023]. In this extension, we explore event classification alongside event detection, expanding our scope to include multi-word event expressions.
@@ 3.1 Source of the short stories
From these eight sources, 25 stories were selected from each for annotation. This rich assortment of stories serves the crucial purpose of delving into a wide range of events and exploring various writing styles and narrative techniques. Each event within these stories is accompanied by an event trigger. These event triggers, typically expressed as verbs, although occasionally as nouns and also adjectives, are specifically designed to elicit the occurrence of that particular event.
@@ 3.2 Annotation Guideline
In this subsection, the key annotation guideline that was utilized for annotation is discussed. Short stories possess unique characteristics that set them apart from other types of texts. Detecting events within short stories poses a specific challenge, especially considering their target audience of children and the narrative structures employed. In these stories, it is common to encounter animals or even inanimate objects engaging in conversations with one another. The occurrence of events in such stories relies heavily on the narrative, genre, and storytelling style. Considering the prior research findings, we have created fresh annotation guidelines centered on recognizing and categorizing event mentions to address these specific challenges for events in short stories. The issues and rules addressed by the guidelines established in ACE Consortium and others [2005], NewsReader Tonelli et al. [2014], HISTO Corpus Sprugnoli and Tonelli [2019], and Litbank Sims et al. [2019a] are taken into account. By leveraging insights from prior research, the aim is to enhance the accuracy and effectiveness of event detection within short stories. The proposed guidelines are valuable resources for annotators, as clear instructions and criteria are provided for identifying and labeling events within the context of children’s short stories. Hereafter, bold words represent realis events, and underlined word shows non-real events that seem to be misunderstood as real events.
@@ 3.2.2 Dimensions to capture realis events
The dimensions followed in the Light ERE approach [Aguilar et al., 2014b] are considered to capture realis events and are mentioned below.
@@
      He traveled to Delhi the last September.
@@ 3.2.3 Event Classes
Every event that has been annotated needs to be categorized to be a particular class. Event categorization is required to know the distribution of different events present in the domain of short stories. As the events in short stories diverge significantly from earlier datasets, adhering to the categorization outlined in prior literature was opted against. The events are classified into seven types so as to cover all the major happenings. The classes are COGNITIVE-MENTAL-STATE (CMS), COMMUNICATION (COM), CONFLICT (CON), GENERAL-ACTIVITY (GA), LIFE-EVENT (LE), MOVEMENT (MOV), and OTHERS (OTH).
@@
Emotional actions like hugging, smiling, laughing, smiling, etc are tagged in this type.
@@
V. LIFE-EVENT: Life event are mostly inspired by the ACE(2005)[Consortium and others, 2005]. These subtypes explore the event tags by LIFE-EVENT.
@@
3.3 Dataset Statistics
The proposed annotation guidelines adapts the problems and rules addressed by ACE [Consortium and others, 2005] for the chosen data. Only the realis events, i.e., the events that are actually happening, with their event classes are considered  [Doddington et al., 2004]. For capturing realis events, four aspects - Polarity, Tense, Genericity, and Modality have been selected [Aguilar et al., 2014a][Song et al., 2015][Sims et al., 2019a].
Sl.No Statistics Count
1. Total tokens in the dataset 174,559
2. Total unique tokens in the dataset 10,484
3. Total sentences in the dataset 10,259
4. Average tokens per story 873
5. Average sentences per story 51
6. Average tokens per sentence 17
7. Total events in the dataset 11,272
8. Average events per story 56
Table 1: Statistics of the annotated dataset (200 short stories)
We annotate 200 stories based on the proposed guideline. An equal number of 25 stories have been considered from each of the sources to avoid biases related to a particular narrative. The web-based BRAT annotation tool [Stenetorp et al., 2012] is used for annotation. The number of words in the stories ranges between 500 to 1500. Two annotators make all the annotations based on the proposed guideline. This study presents an annotated dataset consisting of realis events from Indian short stories. Table 1 shows a detailed breakdown of the dataset statistics. The dataset contains 174,559 tokens, with a significant vocabulary diversity indicated by 10,484 unique tokens. Each story averages 873 tokens and 51 sentences, suggesting moderately lengthy narratives. With an average of 17 tokens per sentence, the stories exhibit concise sentence structures. The dataset records 11,272 events, averaging 56 events per story, indicating a high event density. This suggests that the stories are rich in eventful content, providing a comprehensive resource for studying narrative structures and event annotation.
Event Types Total Count Train Dev Test
COMMUNICATION 3456 2052 447 957
GENERAL-ACTIVITY 2951 1837 325 789
MOVEMENT 2192 1329 294 569
COGNITIVE-MENTAL-STATE 1528 849 178 501
LIFE-EVENT 535 314 60 161
OTHERS 490 295 55 140
CONFLICT 120 62 30 28
Total 11,272 6,738 1,389 3,145
Table 2: Number of different events present in the annotated dataset
Figure 2: Distribution of different events in the dataset
Table 2 categorizes the number of events in the train, development (dev), and test sets across seven event types within the dataset. The total number of events is 11,272, distributed with 6,738 events in the train set, 1,389 in the dev set, and 3,145 in the test set. The most frequent event type is COMMUNICATION, with 3,456 occurrences, followed by GENERAL-ACTIVITY with 2,951 occurrences. MOVEMENT and COGNITIVE-MENTAL-STATE are moderately common, with 2,192 and 1,528 events, respectively. LIFE-EVENT, OTHERS, and CONFLICT are less frequent, indicating they are rarer in the dataset. The distribution across the train, dev, and test sets is consistent, ensuring a balanced representation of each event type, which is crucial for robust model training and evaluation. This detailed annotation provides a rich resource for analyzing and modelling event-based narratives. Figure 2 represents the distribution of the events in the dataset. The dataset clearly shows that COMMUNICATION events are plentiful (30.66%), whereas CONFLICT events are rare (1.06%). The distribution of the events from different short stories is shown in Figure 3. Table 3 represents the top 15 words tagged as an event in the annotated dataset. Figure 4 presents the word cloud of the entire dataset and for trigger words only. It is clear that trigger words with the highest frequencies are in the majority of the entire dataset. This can also be correlated with Table 3. Words like “said”, “asked”, “went”, “came”, “told”, etc., are evident in the dataset.
Two annotators, who are co-authors of this work, performed the annotation task. 20 random stories were chosen, and both the annotators annotated the stories in accordance with the provided annotation guidelines. An inter-annotator agreement (IAA) of 83.1% was achieved. The IAA serves as a meaningful indicator of the clarity of the annotation guidelines, the consistency in the annotators’ comprehension of these guidelines, and the overall feasibility of the annotation task. The remaining 180 stories were annotated by a single annotator, resulting in a total of 200 annotated stories.